# ESGSenticNet: A Neurosymbolic Knowledge Base for Corporate Sustainability Analysis

### v1

@@ 5.5.4 Effectiveness of ESGSenticNet Seed Selection Algorithm
In this baseline study, we assess the extent of label propagation through ESGSenticNet’s seed selection algorithm, against naive methods like selecting seeds randomly or based on highest degree centrality scores. Each method is run on the constructed semantic graph (section 5.4), to select different batches of 10k seeds. Aligning with majority of the classification tasks in the main study, three different psuedo-labels are used, and randomly assigned to the selected seeds. Label propagation is run to observe the number of newly propagated labels. Although propagation would be more accurate with real labels, random pseudo-labelling provides an unbiased estimate of propagation extensiveness irrespective of label content, and reduces the cost of extensive testing. Moreover, the entire process is repeated over 100 trials, showing that across each method, the number of newly propagated labels remain consistent despite the randomness of seed labelling. Compared to naive methods, we observe that ESGSenticNet’s seed selection algorithm results in a greater number of newly propagated labels.
5.6 Seed Annotation
Figure 4: Flowchart depicting inference process
Through GPT-4o few-shot prompting, selected seeds are annotated with their relations (‘aligns with’, ‘supports’, ‘undermines’) with respect to a pillar or category, before their labels are propagated within the semantic graph. Similar to popular sustainability lexicons [6, 54], ESGSenticNet’s approach focuses on seed annotation before leveraging textual similarities to expand the initial seed corpus. Yet, while approaches in the literature annotate a small number of seeds, ESGSenticNet annotates an extensive number of seeds ($>16k$) through GPT-4o. Consequently, ESGSenticNet retains a significant proportion of seed concepts directly labelled through GPT-4o, while non-seeds that rely on textual similarity for labelling are less prevalent. This enriches the diversity of ESGSenticNet, given that the reasoning capabilities of GPT-4o [24] allows it to infer relations for concepts outside of similar semantic spaces. For example, consider the concepts ‘use led light’, ‘achieve energy efficiency’ categorised under ‘Resource Optimisation’. Textual similarity methods leveraging these for comparison might not categorise ‘emphasise water importance’ similarly, despite its relevance. However, using GPT-4o, ‘emphasise water importance’ is appropriately labelled, due to the LLM’s capacity to infer semantic relationships beyond mere textual similarities. Moreover, GPT4o as an automated technique allows for a large number of concepts to be labelled, overcoming prohibitive human labelling costs. This allows us to derive an extensive knowledge base.
To improve GPT-4o classification accuracy, inference tasks are divided into smaller, more manageable sub-tasks [14]. These tasks align with the taxonomy’s hierarchy levels, following the flowchart in figure 4. Subsequent categorisations of seeds are contingent upon their initial classification into broader categories, and it aligns with the predefined associations between the ESG pillars and broad categories, as well as the connections between the broad and sub categories. Prompts for these tasks are highlighted in figure 5 (pillar labelling), figure 6 (relation-topic labelling).
[Task Description]
As a sustainability expert, your task is to classify phrases into the categories of Environmental, Social, Governance, or Others based on their explicit relevance to the Environmental, Social, Governance topics.
[Definitions of Categories]
Others: Reserved for phrases that do not meet the explicitness and specificity requirements for Environmental, Social, or Governance. This includes phrases that are vague, ambiguous, or cover multiple Environmental, Social, Governance topics without clear categorization.
[Response Instructions]
Evaluate each phrase’s explicit content to determine its most appropriate category. Assign each phrase to only one category: Environmental, Social, Governance, or Others. Only phrases that are clear and explicit can be categorized under Environmental, Social, or Governance. Phrases that are vague or not clearly related to these topics should be categorized as Others. Ensure all responses strictly adhere to the definitions and format provided.
[Response Format]
Input: <phrase>
Response: <Environmental, Social, Governance, or Others>
Figure 5: Prompt for labelling an ESG concept’s pillar
[Task Description] As a sustainability expert, you are tasked with evaluating the relationship between phrases and sustainability-related topics. [Definitions of Relationships] supports: The phrase, specific and explicit to the topic, directly advances it with a clear and immediate positive impact. undermines: The phrase, specific and explicit to the topic, directly impedes its goals with a clear and immediate negative impact. [Definitions of Topics] Definitions of topics according to taxonomy [Response Instructions] Evaluate the explicit content of each phrase to determine the most appropriate and accurate relationship and topic. Identify the single most relevant and appropriate relationship and topic. If no specific and explicit relationship can be determined between the phrase and any topic, or if the phrase does not fit any topic due to vagueness or irrelevance, classify the phrase as ‘not applicable’. [Response Format] For each phrase, provide your analysis using the format below. Output only one tuple if applicable, otherwise state ‘not applicable’. Input: <phrase> Response: <(relationship, topic)> or ‘not applicable’
Figure 6: Prompt for labeling an ESG concept’s relation with a topic
5.7 Label Propagation Algorithm
$L^{(k+1)}=D^{-1/2}AD^{-1/2}L^{(k)},$ (7)
$l_{i}^{(k+1)}=l_{i}^{(0)}\quad\forall i\leq m$ (8)
Annotated seeds have their labels propagated to non-seeds via the label propagation algorithm [56]. This increases the labelled data within our knowledge base without incurring additional labelling costs. Specifically, the following algorithm is iterated until convergence, where $A$ is an adjacency matrix containing $a_{i,j}$ that denotes the edge weight between connected nodes $e_{i}$ and $e_{j}$. Each node represents an embedding of a concept, while the edge weight denotes cosine similarity between concepts. $L$ is a matrix that represents the relations labels of each node with respect to a topic (i.e. supports within the (‘reduce water consumption’, supports, ‘resource optimisation’). $m$ is the number of nodes possessing labels and $D$ is a diagonal matrix containing $d_{i,i}$, which equals the sum of elements in row $i$ of $A$. We use parameters $n\_layers$=50, $\alpha$=0.5.
5.8 ESGSenticNet Statistics
Categories Total Triplets Number of triplets with the relations

### v2

@@ 4.10.1 Effectiveness of ESGSenticNet Seed Selection Algorithm
In this baseline study, we assess the extent of label propagation through ESGSenticNet’s seed selection algorithm, against naive methods like selecting seeds randomly or based on highest degree centrality scores. Each method is run on the constructed semantic graph (Section 4.8.1), to select different batches of 10k seeds. Aligning with majority of the classification tasks in the main study, three different psuedo-labels are used, and randomly assigned to the selected seeds. Label propagation is run to observe the number of newly propagated labels. Although propagation would be more accurate with real labels, random pseudo-labelling provides an unbiased estimate of propagation extensiveness irrespective of label content, and reduces the cost of extensive testing. Moreover, the entire process is repeated over 100 trials, showing that across each method, the number of newly propagated labels remain consistent despite the randomness of seed labelling.
@@
Compared to naive methods, from Table 11, we observe that ESGSenticNet’s seed selection algorithm results in a greater number of newly propagated labels.
4.10.2 GPT-4 Human Expert Agreement on Topic Analysis Evaluation Metrics
GPT-4 evaluations have shown significant reliability in the literature [80]. In line with established research practices [61], we validate the GPT-4 evaluations in our use case by conducting a GPT4-human agreement study. Specifically, a study is conducted to assess GPT-4’s ability for the following tasks– i) classifying topic terms as ESG-related, ii) classifying topic terms as actions taken toward ESG. A random sample of 510 topic terms were gathered for each task. GPT-4 annotations for each set of topic terms are compared with annotations from 2 different groups of 3 human experts each, with each group handling a different task. These human experts comprise of doctoral candidates actively conducting research in sustainability finance, and thus possess domain expertise well beyond the level required for this validation task. Moreover, to ensure the robustness of their judgements, the human experts were deliberately kept independent from the paper’s development. An average agreement score of 83.7% observed for task i) and 91.8% for task ii). While this highlights a reasonable level of robustness for GPT-4 evaluations in our study, it also underscores the broader potential of GPT-4 as a scalable and reproducible evaluation tool in the ESG domain.
4.10.3 Effectiveness of ESGSenticNet Flexible Matching Method
A study is conducted to assess the precision of matching ESGSenticNet concepts through the ‘flexible’ method depicted in Section 2.2. A random sample of 200 sentences matched through this method was collected, excluding sentences matched via the ‘precise’ method. Human annotators were tasked with evaluating whether each sentence clearly conveys the idea expressed represented by its matched concept. From our results, 75.0% of sentences comprised the idea expressed by its corresponding matched concepts. This suggests that the ‘flexible’ method retains a moderately high level of precision, despite allowing for greater semantic variability. For evaluating the ESGSenticNet flexible matching method, the following instructions were given to evaluate if each sentence clearly conveys the idea expressed by its matched concept. The table below presents different sentences and a corresponding concept. Determine if each sentence clearly conveys the idea expressed by its matched concept. Mark ‘TRUE’ if the sentence clearly conveys the concept’s idea, and ‘FALSE’ otherwise.
5 Literature Review
5.1 NLP-driven ESG topic analysis.
To extract insights from sustainability reports, NLP tools can be deployed for different tasks, such as ESG topic analysis [28, 41], sentiment classification [58], extracting textual patterns [9], with ESG topic analysis being a key task [54]. ESG topic analysis facilitates ESG assessments in several ways. Specifically, analysing topics allows stakeholders to monitor corporate alignment with UN Sustainability Development Goals [70], to analyse the future orientation of sustainability reports [34], and to construct metrics with respect to ESG performance [68].
Current NLP methods for this task predominantly include rule-based lexical methods that leverage existing sustainability dictionaries such as Baier [37], and unsupervised topic modeling that can be adapted to ESG analysis [82, 45]. Yet, while having advantages as automated tools [2], these methods may not be tailored for the specialised requirements of ESG topic analysis. Moreover, building specific tools for ESG topic analysis has been under-explored, with efforts primarily focusing around the creation of sustainability lexicons [5, 60, 39]. However, these lexicons have notable limitations, which we will discuss further in the next section.
5.2 Lexicons for ESG Analysis
Lexical approaches have been widely used to analyse sustainability reports [34], given their effectiveness in corporate text analysis and accessibility to stakeholders with diverse technical backgrounds [51, 11]. To facilitate lexical analysis, lexicons specific to ESG analysis have been developed. For instance, approaches such as Baier’s dictionary [5], have focused on developing single-word lexicons according to ESG pillars. However, while Baier’s dictionary has been a popular choice for ESG analysis studies [37], its single-word lexicons cannot adequately capture the semantics of the sustainability field [65]. In contrast, dictionaries that have collected ESG-related text spans (of multiple-grams)—i.e.  Du et al. [23]Kang and El Maarouf [39]—possess limited ESG topics or are not publicly available. Additionally, due to their reliance on manual annotation, all the aforementioned dictionaries have an exceedingly small number of lexicons (less than 800). This raises questions on their comprehensiveness and consequently their effectiveness for capturing textual patterns. To address these limitations, our constructed knowledge base introduces a structured lexicon of ESG concepts—one that includes multiword expressions that more effectively capture the semantics of the sustainability field, and is more extensive, comprising 23k concepts.
Additionally, existing publicly available ESG lexicons—which, to the best of our knowledge, predominantly include Baier [5] and Naiara [60]—represent structured attempts to derive ESG knowledge, but differ significantly from our neuro-symbolic framework in both methodology and scalability. Baier’s dictionary is constructed through manual annotation and term frequency heuristics applied to sustainability documents, limiting its scalability and adaptability. Naiara’s dictionary, by contrast, derives ESG-related terms via cosine similarity to a small set of predefined anchors (“environmental,” “social,” and “governance”), which restricts the diversity of retrieved concepts and biases term selection toward those semantically proximal to the anchors. In contrast, our approach is fully automated and designed to support both scalability and conceptual diversity. It combines GPT-4o-based annotation with label propagation, uses automated concept parsing to extract candidate ESG terms, and applies data clustering and graph-based analysis to select seed terms from semantically diverse regions of the data space. Compared to existing lexicons, our framework yields a relatively more diverse and extensive repository of ESG lexicons (or concepts), organised within a structured topical hierarchy. As highlighted by our experimental results in the later section (Section 2.3), this provides a greater number of unique ESG terms extracted in ESG topic analysis.
5.3 ESG Knowledge Graphs
While prior works have applied knowledge graphs to the analysis of sustainability data [31, 4], these efforts are primarily constructed to support downstream tasks such as relation extraction or question answering over ESG-related news or sustainability reports. In these cases, the knowledge graphs are designed to represent document-specific entities and relations, but are not formalized as reusable or publicly accessible repositories of ESG concepts. Moreover, these approaches are not oriented toward ESG topic analysis—they do not aim to identify, structure, or surface ESG topics and topical terms systematically across corpora. By contrast, ESGSenticNet is explicitly designed as a structured and reusable knowledge base of ESG concepts: it organizes ESG terms and their topical relations into a coherent taxonomy, and is openly available for external use. While ESGSenticNet can also support downstream applications such as relation extraction, its foundational role is to serve as a general-purpose, publicly accessible repository for ESG topic analysis. Furthermore, because ESGSenticNet can be deployed through a structured lexical method, it requires no training or technical expertise—making it accessible to stakeholders across diverse backgrounds.
5.4 Semantic Knowledge Bases
In the literature, semantic knowledge bases are often represented as knowledge graphs consisting of triples in the form of (subject, predicate, object), or (head, relation, tail), which encode structural relations between entities [38]. By structuring knowledge in a way that mimics human cognition, knowledge bases have been explored in numerous NLP applications, including commonsense question answering [20], sarcasm detection [78], and sentiment analysis [81], among others, as part of broader research efforts toward more context-aware and human-like language understanding. However, the construction of knowledge bases within the ESG domain remains largely unexplored. To address this gap, this paper presents the first attempt at developing a knowledge base that organises ESG knowledge into structured relationships.
Methods for constructing knowledge bases can vary according to the domain and application [38]. Notable knowledge bases, such as ConceptNet [42] involved concept parsing and human annotation, SenticNet [14] leveraged deep learning and lexical substitution, other knowledge bases involve question answering via advanced graphical neural networks (GNNs) [77]. Newer avenues for knowledge construction involve generative-LLMs, which have shown strong performance in related tasks such as reasoning, construction, and generalisation [83]. Unlike all the previous methods which heavily rely on semantic similarities for labelling, generative LLMs possess powerful emergent capabilities that extend beyond textual similarities [72]. These include reasoning, knowledge relationship construction, and generalisation to different domains [83]. Moreover, unlike GNNs that rely on heavily structured text data [77], generative-LLMs can flexibly adapt to unstructured text [7]. Yet, leveraging generative LLMs for knowledge construction is relatively new, with limited exploration of its effectiveness. This study contributes to this area by constructing a ESG-focused knowledge base using generative LLMs, and evaluating the resulting knowledge base’s accuracy in categorising ESG concepts under the correct topics.
5.5 Neurosymbolic Methodologies
Neurosymbolic approaches integrate symbolic AI (rule-based systems) with sub-symbolic AI (neural networks) to enhance generalisation, interpretability, and reasoning in AI [33]. These methods have been successfully applied to various NLP tasks, including word sense disambiguation for sentiment analysis [79], recommender systems [66], and question-answering [20]. However, their potential for large-scale knowledge base construction remains largely unexplored. This study extends neurosymbolic methodologies by developing a neurosymbolic framework for knowledge construction. Specifically, sustainability linguistic patterns are integrated into rule-based concept extraction while GPT-4o is used for labelling concepts’ relations with ESG topics. In doing so, this study explores the potential of a neurosymbolic approach to enhance the relevance and accuracy of knowledge construction.
6 Conclusion
In this study, we described the challenges of ESG topic analysis, outlining the limitations of existing NLP approaches—i.e., irrelevance, immateriality, limited organization. To specifically address these challenges, we developed a knowledge base, ESGSenticNet, from a novel neurosymbolic framework, before evaluating its effectiveness against existing NLP benchmarks. The results indicate that ESGSenticNet captures more relevant, material, and comprehensive insights for ESG topic analysis, significantly outperforming existing sustainability dictionaries and state-of-the-art NLP methods. Additionally, the topic terms ESGSenticNet provides are correctly associated with their intended topics to a reliable degree, highlighting its robustness for topic analysis. All in all, ESGSenticNet’s deployment as a structured lexicon method does not require computational training or technical expertise, meaning it can be readily used by stakeholders regardless of their technical background. Ultimately, we hope that ESGSenticNet democratises access to crucial insights for ESG assessments.
7 Limitations
Although ESGSenticNet is designed to be an accessible and inclusive tool for ESG topic analysis, it has several limitations. First, due to the evolving nature of ESG frameworks [54], ESGSenticNet may require periodic updates to incorporate newly emerging ESG topics. Second, the resource is currently limited to concepts expressed in English and is not yet applicable to non-English sustainability reports. Future work will therefore focus on extending ESGSenticNet to multiple languages, potentially through the use of multilingual large language models [50]. Third, while ESGSenticNet covers a broad range of ESG topics and is built on a large corpus of sustainability reports, it does not explicitly assess the generalisability of its concepts across reports from different industries (e.g., finance vs. energy). In other words, our study does not investigate whether ESGSenticNet provides equally effective topic analysis for sustainability reports from companies in different sectors. Addressing this may involve developing cross-industry generalisation tasks and exploring learning algorithms (i.e. contrastive learning) to improve robustness [55]
Additionally, while our neurosymbolic approach mitigates overreliance on LLMs (specifically GPT-4o) by incorporating sustainability-specific linguistic patterns through symbolic concept parsing, we acknowledge that LLMs may still exhibit category-specific biases due to their pretraining [21]. For instance, they may be more accurate in classifying relations for certain sustainability categories over others [55]. To address this, we decompose the LLM inference task into smaller, focused sub-tasks (Section 4.3.3), improving accuracy even with less familiar categories [17]. We further validate the reliability of LLM classifications through a human evaluation study detailed in Section 5.4.
Declarations
The authors declare that they have no competing interests as defined by Nature Portfolio, or other interests that might be perceived to influence the results and/or discussion reported in this paper. Author Contributions: K. O. : Writing – review & editing, Writing – original draft, Methodology, Investigation, Formal analysis, Conceptualisation. R, M.: Writing – review & editing, Methodology, Investigation, Concep- tualisation. D.V.: Conceptuslisation, F.X: Conceptualisation. R. S.: Resources. J. S.: Writing – review & editing, Methodology, Investigation. E. C. Writing – review & edit- ing, Conceptualisation, Supervision. G. M.: Writing – review & editing, Supervision, Methodology, Conceptualisation.
Acknowledgements & Funding Statement
This research/project is supported by the NUS Sustainable and Green Finance Institute (SGFIN), NUS Asian Institute of Digital Finance (AIDF), Ministry of Education, Singapore under its MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005), MOE Tier 2 Award (MOE-T2EP50221-0006: “Prediction-to-Mitigation with Digital Twins of the Earth’s Weather”), MOE Tier 1 Award (MOE-T2EP50221-0028: “Discipline-Informed Neural Networks for In- terpretable Time-Series Discovery”), and by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore.
Appendix A ESGSenticNet Taxonomy
Table 16 contains the definitions of the pillars (Environmental, Social, Governance) and topics used in ESGSenticNet. Along with the pillars, the broad topics, and cross-broad topics are bolded, with their constituent sub-topics highlighted in bullet points. Beside the broad and cross-broad topic names, a letter denotes which pillar the topic falls under – (E) Environmental, (S) Social, (G) Governance.
Topic Description
Table 12: ESGSenticNet Sustainability Taxonomy - continued in the next page
Table 13: ESGSenticNet Sustainability Taxonomy - continued in the next page
Table 14: ESGSenticNet Sustainability Taxonomy - continued in the next page
Table 15: ESGSenticNet Sustainability Taxonomy - continued in the next page
Table 16: ESGSenticNet Sustainability Taxonomy
Appendix B Additional details on Human Validation
For evaluating ESGSenticNet accuracy (Section 2.4), ESGSenticNet’s flexible matching method (Section 4.10.3), GPT-4 human agreement (Section 4.10.2), there are a total of 9 human annotators. These annotators are based in Singapore, comprising researchers from the Asian Institute of Digital Finance and the Sustainable and Green Finance Institute, who are pursuing doctoral-level and post-doctoral research within the corporate sustainability field. All human annotators contribute to this study as part of their formal academic and research responsibilities. All human validation studies follow the same annotation scheme. When there are disagreements, discussions are held to achieve a resolution. Where there is no clear resolution, majority voting is taken to determine the ground truth label. For the derivation of the taxonomy for ESGSenticNet, we involved two tenured researchers from the Singapore Sustainable and Green Finance Institute (SGFIN), with expertise within the sustainable finance field.
Appendix C Full Statistical Breakdown of ESGSenticNet
Topics Total triples Number of triples with the relations